OpenStudy (anonymous):

simplify sin(3x)/tan(3x)

OpenStudy (anonymous):

Help? I'm stuck big time...

OpenStudy (amistre64):

what is another way of writting 1/tan ?

OpenStudy (anonymous):

Cos/sin so then I should get cos 3x, right?

OpenStudy (anonymous):

Hint:\[\large \tan 3x = \frac{\sin 3x}{\cos 3x}\]

OpenStudy (anonymous):

Correct!

OpenStudy (anonymous):

Aha! Thanks so much you two!! I've got it from here!